OpenStudy (ccchristian1):

tell whether the lines for each pair of equations are parallel perpendicular, or neither. y=-4x+3 -2x+8y=5

OpenStudy (anthonyym):

I would put the second equation in slope intercept form y=mx+b first.

OpenStudy (ccchristian1):

@ivettef365

OpenStudy (anthonyym):

In other words, solve -2x+8y=5 for y.

OpenStudy (ccchristian1):

i got a

OpenStudy (ccchristian1):

idk if im right or wrong

OpenStudy (ivettef365):

find the slope of each equation

OpenStudy (ivettef365):

to be parallel the slope have to be the same to be perpendicular the slope have to be opposite reciprocal otherwise is neither

OpenStudy (ivettef365):

on the first equation you have y = -4x + 3 so what is the slope ?

OpenStudy (ccchristian1):

how would u get neither?

OpenStudy (ccchristian1):

-4 is the slope

OpenStudy (ivettef365):

if both equation have the same slope is parallel if one equation have a slope that is opposite reciprocal from the other is perpendicular and what I mean by opposite reciprocal if one have a slope of -4 and the other 1/4 that is opposite reciprocal

OpenStudy (ivettef365):

if the slope is NOT the same and is NOT opposite reciprocal then is neither

OpenStudy (ccchristian1):

so -4 and 4 wouldn't be opposite ?

OpenStudy (ivettef365):

yes that would be just opposite

OpenStudy (ccchristian1):

alright i get it now

OpenStudy (ccchristian1):

i wish my teachers would explain it how u explain it

OpenStudy (ivettef365):

but to be opposite reciprocal no only change from pos to neg or neg to opp also the number switch for example if you have 3/4 then opp rec would be -4/3

OpenStudy (ccchristian1):

ahh okay

OpenStudy (ivettef365):

so now you have y - -4x +3 you know in the formula y = mx + b the number of the position of the m is the slope

OpenStudy (ccchristian1):

so its b?

OpenStudy (ivettef365):

the slope is the m the b is the y-intercept

OpenStudy (ccchristian1):

thats my answer rn

OpenStudy (ivettef365):

it is perpendicular because the second equation is -2x + 8y = 5 8y = 2x + 5 y = 1/4x + 5

OpenStudy (ccchristian1):

alright so i was right thank u for explaining it to me

OpenStudy (ivettef365):

the first slope is -4 and the second slope is 1/4 that is opposite and reciprocal

OpenStudy (ivettef365):

you're welcome

OpenStudy (ccchristian1):

1/4 and -1/4 would be perpendicular.. correct ?

OpenStudy (ccchristian1):

@ivettef365

OpenStudy (ivettef365):

yes

OpenStudy (ccchristian1):

okay thank you

OpenStudy (ivettef365):

no problem anytime :)